OpenStudy (anonymous):

solve and check equation: (x^2+16x+64)^3/4-12=15

OpenStudy (anonymous):

x=-17, x=1

OpenStudy (anonymous):

how did you get that?!

OpenStudy (anonymous):

It's a quadratic equation. Set equal to zero. Solve by combining like terms.

OpenStudy (anonymous):

If you need more information give me a second. lol

OpenStudy (anonymous):

okay thanks ill wait :)

OpenStudy (anonymous):

okay thanks ill wait :)

OpenStudy (anonymous):

(x^(2)+16x+64)^((3)/(4))-12=15 Since -12 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 12 to both sides. (x^(2)+16x+64)^((3)/(4))=12+15 Add 15 to 12 to get 27. (x^(2)+16x+64)^((3)/(4))=27 Eliminate the fractional exponent on the lefthand side by raising both sides by to the (4)/(3) power. ((x^(2)+16x+64)^((3)/(4)))^((4)/(3))=(27)^((4)/(3)) Expand the exponent of (4)/(3) to each factor in the expression (x^(2)+16x+64)^((3)/(4)). ((x^(2)+16x+64))=(27)^((4)/(3)) Remove the parentheses from the numerator. (x^(2)+16x+64)=(27)^((4)/(3)) Remove the parentheses around the expression x^(2)+16x+64. x^(2)+16x+64=(27)^((4)/(3)) Expand the exponent ((4)/(3)) to the expression. x^(2)+16x+64=(27^((4)/(3))) An expression with a fractional exponent can be written as a radical with an index equal to the denominator of the exponent. x^(2)+16x+64=((~3:(27))^(4)) Pull all perfect cube roots out from under the radical. In this case, remove the 3 because it is a perfect cube. x^(2)+16x+64=((3)^(4)) Expand the exponent (4) to the expression. x^(2)+16x+64=((3^(4))) Raising a number to the 4th power is the same as multiplying the number by itself 4 times. In this case, 3 raised to the 4th power is 81. x^(2)+16x+64=((81)) Remove the parentheses from the numerator. x^(2)+16x+64=(81) Remove the parentheses from the numerator. x^(2)+16x+64=81 To set the left-hand side of the equation equal to 0, move all the expressions to the left-hand side. x^(2)+16x-17=0 In this problem 17*-1=-17 and 17-1=16, so insert 17 as the right hand term of one factor and -1 as the right-hand term of the other factor. (x+17)(x-1)=0 Set each of the factors of the left-hand side of the equation equal to 0. x+17=0_x-1=0 Since 17 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 17 from both sides. x=-17_x-1=0 Set each of the factors of the left-hand side of the equation equal to 0. x=-17_x-1=0 Since -1 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 1 to both sides. x=-17_x=1 The complete solution is the set of the individual solutions. x=-17,1

OpenStudy (anonymous):

too much? lol

OpenStudy (anonymous):

thank you soo much but you could have just written it out! :p

OpenStudy (anonymous):

haha, well it's important to understand whats confusing you. I have a program that generates everything if someone needs details. : ) I originally solved on paper, but thought a detailed answer might help you more!

OpenStudy (anonymous):

aha yes thank you though!! i mean i had the basics i just didnt equal everything to 0 i would put 17 there instead or w.e

OpenStudy (anonymous):

yeah, just remember, if you have something with x^2 in it and x then a constant (a number) there's a good chance you can use quadratics, or even the quadratic formula to solve it!